OpenStudy (anonymous):

when n is divided by 14, the remainder is 10. what is the remainder when n is divided by 7?

OpenStudy (gw2011):

The answer is a remainder of 3. Here's how I got the answer: First, by definition the dividend is the number that is being divided; the divisor is the number that is dividing the dividend; and the quotient is the result. It is a fact that: Dividend=(Divisor)(Quotient)+Rmainder For example: 19/5=3 R4 where 19=Dividend, 5=Divisor, 3=Quotient, and 4=Remainder. Since we know that: Dividend=(Divisor)(Quotient)+Remainder, we get: Dividend=(5)(3)+4=15+4=19 We can now apply this to the problem using the following formula: Dividend=(Divisor)(Quotient)+Remainder or using symbols to get: N=(D)(Q)+R where N=Dividend where D=Divisor where Q=Quotient where R=Remainder For N/14, we get: N=(14)(Q)+10 Since Q (the Quotient) has to be a whole number because of what the problem states, Q can equal 1 or 2 or 3 or 4, etc. If Q=1: N/14: N=(14)(1)+10=24 24/14=1 R10 N/7=24/7 (the number 24 is from above)=3 R3 If Q=2: N/14: N=(14)(2)+10=28+10=38 38/14=2 R10 N/7=38/7=5 R3 Therefore, in each case the remainders are the same: for N/14, the remainder=10, and for N/7, the remainder=3. Consequently, the formula works and the remainder for N/7 is 3.